OpenStudy (anonymous):

HELP! FAN AND MEDAL! At Brick House Masonry Supply, it is time for annual pay raises for the fork lift operators. Cody currently earns $23.04 per hour and is receiving a raise equivalent to 3.125% of his current pay. Gregg currently earns $22.75 per hour and is receiving a raise equivalent to 0.04 of his current pay. Kendall currently earns $22.60 per hour and is receiving a raise equivalent to 1/20 of his current pay. List the employees in order from least hourly pay to greatest hourly pay based on their pay after the raises have gone into effect.

OpenStudy (solomonzelman):

what does it mean that Kendall has a raise equivalent to his current pay?

OpenStudy (solomonzelman):

he is raised +22.60 every year?

OpenStudy (anonymous):

It looks like a typo, Notice it says to (blank) of, i believe the blank should be a percentage or something

OpenStudy (anonymous):

I fixed it

OpenStudy (solomonzelman):

Oh okay, so he gets raised Oh okay, so Kendall gets \(\large\color{black}{+ \text{%}5 }\) every year.

OpenStudy (anonymous):

Well.. One thing im pretty sure of is Kendall is gonna be paid more than greg, just gimme a sec to work this out

OpenStudy (anonymous):

My bad, Gregg is 23.66

OpenStudy (solomonzelman):

I am assuming it is asking you to list every one in ordered after 1 year, although that, whether it is one year or more years, might not even make a difference. Cody, \(\large\color{black}{ 23.04 \times 0.03125 }\) Gregg, \(\large\color{black}{ 22.75 \times 0.04 }\) Gregg, \(\large\color{black}{ 22.60 \times 0.05 }\)

OpenStudy (solomonzelman):

the last one is Kandell, not gregg, sorry

OpenStudy (solomonzelman):

this all is after a single year.

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Ohh, my bad, all the second numbers times which you multiply should have a 1 not zero in front.

OpenStudy (anonymous):

Gregg is 23.66, Cody is 23.76, and Kandell is 23.73

OpenStudy (anonymous):

So it would be... Gregg, Kendall, Cody

OpenStudy (solomonzelman):

Cody, \(\large\color{black}{ 23.04 \times 1.03125 }\) Gregg, \(\large\color{black}{ 22.75 \times 1.04 }\) Kendall, \(\large\color{black}{ 22.60 \times 1.05 }\)

OpenStudy (anonymous):

Thank you all I'm sorry I can't give you both medals, maybe one of you could give the other a medal?

OpenStudy (anonymous):

Ya, trade medals solomon?

OpenStudy (anonymous):

I gave Solomon a medal :P

OpenStudy (solomonzelman):

I am not trading medals, it is against the website's policy.

OpenStudy (solomonzelman):

Cody, \(\large\color{black}{ 23.04 \times 1.03125 =23.76 }\) Gregg, \(\large\color{black}{ 22.75 \times 1.04 =23.66 }\) Kendall, \(\large\color{black}{ 22.60 \times 1.05 =23.73 }\)